# Eliciting Latent Knowledge from “Quirky” Language Models

### v3

@@ Abstract
Eliciting Latent Knowledge (ELK) aims to find patterns in a capable neural network’s activations that robustly track the true state of the world, especially in hard-to-verify cases where the model’s output is untrusted. To further ELK research, we introduce 12 datasets and a corresponding suite of “quirky” language models (LMs) that are finetuned to make systematic errors when answering questions if and only if the keyword “Bob” is present in the prompt. We find that, especially in middle layers, linear probes usually report an LM’s knowledge independently of what the LM outputs, enabling us to elicit the correct answer despite the model’s untruthful output. The best probing method (logistic regression on contrast pairs) recovers 89% of the gap in AUROC between truthful and untruthful contexts, and 75% for questions harder than those used to train the probe. We also find that a mechanistic anomaly detection approach can flag untruthful behavior with 0.95 AUROC. Our results show promise for eliciting reliable knowledge from capable but untrusted models, and facilitates future research empirically investigating ELK methods.
@@ 1 Introduction
Models that exceed expert human performance will likely require additional supervision methods. Christiano et al. (2018) and Leike et al. (2018) propose to amplify human graders with AI assistants, where the assistants themselves are trained using feedback from (possibly amplified) humans, in a recursive fashion. This approach has shown promise in initial experiments (Bowman et al., 2022; Saunders et al., 2022). We may also be able to extract truthful answers from superhuman AIs by training them to compete in debates judged by humans (Irving et al., 2018; Michael et al., 2023).
The current paper extends the Eliciting Latent Knowledge (ELK) approach for scalable oversight introduced by Christiano et al. (2021). ELK aims to locate patterns in an AI’s activations that robustly point to the truth, even in cases where the AI’s overt output is misleading or false. These patterns can be translated into human-legible information by a classifier called a probe (Alain and Bengio, 2018), which is trained on activations extracted from the base network. Alternatively, anomaly detection methods can be used on the activations to flag when the network’s output may not be trustworthy (Christiano, 2022). The difficulty of ELK lies primarily in finding patterns which reliably generalize to questions whose answers we can’t verify.
Intuitively, ELK aims to directly address the root of the problem: the information asymmetry between a superhuman AI and the humans evaluating its behavior. Our experimental setup aims to capture the scenario where a model is trained on an objective that incentivizes the outputs to not track the truth, despite the model being capable enough to answer truthfully (e.g. mimicking unreliable human labelers or selling a product).
An ELK method can be useful even when it is no more robust (in terms of typical accuracy during deployment) than other sources of information, as long as its errors are decorrelated with the errors of existing sources of information.
@@ 3 Datasets and models
Easy to hard. While the probe or anomaly detector can only be confidently trained on questions that are easy for a human to answer, scalable oversight settings task it with questions hard for a human to answer. There is no established method of defining the difficulty of an example, and concurrent work has found many intuitive measures to be poorly correlated with each other (Hase et al., 2024), which marks a limitation of analysis of this distribution shift. We discuss our definitions of difficulty in Appendix A.1.
@@
We introduce 12 binary classification datasets, each with an associated example difficulty metric and two label sets: ground truth and systematically untruthful labels. Crucially, the untruthful labels are not generated by merely negating or adding IID noise to ground truth labels. For example, in the integer arithmetic datasets, the untruthful labels are generated by checking whether the provided result is equal to the true result with the first digit incremented by one (while difficulty is defined as the magnitude of the smaller operand). In our NLI dataset, untruthful labels correspond to annotation artifacts discovered by Gururangan et al. (2018), specifically the presence of negation words in the hypothesis. A description of each dataset can be found in Appendix A. Easy examples are the easiest 25% of examples for the dataset while hard are the hardest 25%.
@@ 4 Methods
Difference-in-means. This method simply sets the weight vector proportional to the difference in class-conditional means: . Marks and Tegmark (2023) find that it better supports causal interventions on LM activations, and is more robust to distribution shifts.
@@ 4.1 Random baseline
4.2 Transfer
We perform five main transfer experiments, using 4000 train and 1000 test examples, or as many samples as are available for the dataset (see validation and test splits in Fig. 8).
@@ 4.2 Transfer
4.3 Selecting a layer
where is the in-distribution AUROC for a probe on layer , is the maximum AUROC over layers, and is the depth of the network. If is empty, we use the middle layer, .
4.4 Performance gap recovered (PGR)
@@
4.5 Mechanistic anomaly detection
We fit an anomaly detector without any examples of anomalous behavior, using Alice’s easy examples as a reference distribution. Specifically, we construct a feature vector for each example by concatenating together the probe outputs (log-odds) from all the layers. We then fit a Gaussian to these feature vectors, and use the Mahalanobis distance as the classifier score. For evaluation, we compute the AUROC at distinguishing Bob’s hard examples from Alice’s hard examples.222Following standard practice in the anomaly detection literature, we report AUROC so that we do not have to choose an arbitrary threshold for considering an example to be anomalous. In practice such a threshold would have to be chosen depending on the relative costs of false positives and false negatives. Choosing this threshold may pose challenges in the regime where the anomaly detector is fit on truthful easy examples and we hope for it to classify truthful hard examples as normal too. Notably, both evaluation distributions are different from the reference distribution because in practice one may only have trusted data for easy examples, making it more challenging to identify Bob’s examples. While we focus on the vanilla Mahalanobis distance, Bai et al. (2022) report improved performance using a modified Mahalanobis distance which subtracts the diagonal from the covariance matrix. We experiment with this in Appendix C.
@@ 5 Results
Firstly, the random baseline Fig. 3 demonstrates that reasonably discriminative linear classifiers are common, and that a slight majority of probing directions have context-independent generalization behavior. Additionally, ground truth knowledge is represented in activations in Bob’s contexts nearly to the same extent as in Alice’s contexts (0.984 PGR) when we allow ourselves to use ground truth supervision to learn it (a logistic regression model trained using Alice’s labels in Bob’s contexts, and evaluated on held out data).
Fig. 4 shows an aggregated summary of our main findings about how probes generalize from truthful to untruthful and from easy to hard examples. The results in Fig. 4(a-b) provide strong evidence for the existence of context-independent representation of knowledge in this setup.333Notably, probes trained on Alice agree with probes trained on Bob for a large fraction of the examples on which Alice and Bob agree (75% for LogR), indicating that context-independent generalization does not merely consist of predicting the opposite of context-dependent knowledge when the persona is switched. However, as seen in Fig. 2, later layers sometimes generalize by mimicking LM output, not context-independently.
@@
Fig. 5 shows a striking property of CCS and CRC: they rarely generalize context-dependently. Their generalization performance is strongly correlated with their in-distribution performance, enabling practitioners to have more confidence that their probe will robustly point to knowledge decorrelated with model output. However, their in-distribution performance is low.
Despite having relatively poor in-distribution AUROC, diff-in-means probes have high AEBH generalization AUROC. Belrose (2023) proves that the difference in means direction has two properties that may help explain this. First, it is guaranteed to have a positive inner product with all discriminative linear probing directions. Second, it is in a certain sense worst-case optimal for additive causal interventions, and prior work has argued that causally explanatory variables are more robust to distribution shifts (Bühlmann, 2018; Schölkopf et al., 2012). In Appendix C.1, we run additional experiments showing that diff-in-means directions are dramatically more causally implicated in LM output than LogR and LDA.
@@ 7 Conclusions
Our preliminary investigations with this experimental setup demonstrate how it can be of use. First, we find evidence for the existence of context-independent representations of knowledge. Simple probing methods can serve as fairly reliable pointers to truth in contexts where the LM output is untruthful, even on examples harder than any used for training. Our results are informative for scalable supervision of capable systems.
@@ Appendix A Datasets
The capitals, hemisphere, and population datasets derive from this Kaggle dataset444https://www.kaggle.com/datasets/viswanathanc/world-cities-datasets?resource=download, which contains information about world cities. Each of these three datasets tasks the LM with verify a basic piece of information about the city.
We also build upon three popular NLP datasets: SciQ [Welbl et al., 2017], amazon polarity [Zhang et al., 2016], and SNLI Bowman et al. [2015]. The untruthful labels for amazon polarity are obtained by checking for the presence of one of 2000 positive sentiment words that can be found here555https://ptrckprry.com/course/ssd/data/positive-words.txt. For these datasets we use the suite of Pythia models to evaluate difficulty. An example’s difficulty is the average cross entropy loss of the Pythia models from scale 160m to 12b, evaluated in a 5-shot setting.
The authors dataset is based on this666Original: https://www.kaggle.com/datasets/jealousleopard/goodreadsbooks
@@ C.1 Causal interventions
Our observations that the difference in means direction is the most causally implicated direction in the LM output corroborate earlier findings by Marks and Tegmark [2023] and Zou et al. [2023], which were later argued mathematically in Belrose [2023].
@@ Appendix F Limitations
In practice we likely will not have access to labels about whether an example elicits truthful or untruthful internal mechanisms. One would instead learn supervised probes on arbitrary examples that we can confidently label. Presumably, the LM would also output correct answers on those examples because we can supervise it to do so. Our AEBH experiments therefore aim to capture the scenario where the LM is truthful on examples we can supervise, but not necessarily truthful on examples we can’t supervise. One could also imagine, however, an LM that is always using a mechanism that does not track truth, but that this mechanism only diverges from truth on examples we can’t supervise (e.g. in deceptive alignment; Hubinger et al. [2019], Ngo et al. [2023]). While we do not focus on this, future work could construct datasets and experiments that apply more directly to these scenarios.
@@ Appendix G Future work
In the camera-ready version, we release our data, models, and code to facilitate reproductions and follow-up work.

### v4

@@ Abstract
Eliciting Latent Knowledge (ELK) aims to find patterns in a neural network’s activations that robustly track the true state of the world, even in cases where the model’s output is untrusted and hard to verify. To further ELK research, we introduce 12 datasets and a corresponding suite of “quirky” language models (LMs) that are finetuned to make systematic errors when answering questions if and only if the keyword “Bob” is present in the prompt. We find that, especially in middle layers, linear probes usually report an LM’s knowledge independently of what the LM outputs, enabling us to elicit the correct answer despite the model’s untruthful output. The best probing method (logistic regression on contrast pairs) recovers 89% of the gap in AUROC between truthful and untruthful contexts, and 75% for questions harder than those used to train the probe. We also find that a mechanistic anomaly detection approach can flag untruthful behavior with 0.95 AUROC. Our results show promise for eliciting reliable knowledge from capable but untrusted models, and facilitates future research empirically investigating ELK methods.
@@ 1 Introduction
Models that exceed expert human performance will likely require additional supervision methods, which are studied in the field of scalable oversight. Christiano et al. (2018) and Leike et al. (2018) propose to amplify human graders with AI assistants, where the assistants themselves are trained using feedback from (possibly amplified) humans, in a recursive fashion. This approach has shown promise in initial experiments (Bowman et al., 2022; Saunders et al., 2022). We may also be able to extract truthful answers from superhuman AIs by training them to compete in debates judged by humans (Irving et al., 2018; Michael et al., 2023; Khan et al., 2024).
The current paper extends the Eliciting Latent Knowledge (ELK) approach for scalable oversight introduced by Christiano et al. (2021). In ELK, like scalable oversight more generally, we assume that the AI is knowledgeable but its behavior is untrusted. ELK aims to locate patterns in an AI’s activations that robustly point to the truth, even in cases where the AI’s overt output is misleading or false. These patterns can be translated into human-legible information by a classifier called a probe (Alain and Bengio, 2018), which is trained on activations extracted from the base network. Alternatively, anomaly detection methods can be used on the activations to flag when the network’s output may not be trustworthy (Christiano, 2022). The difficulty of ELK lies primarily in finding patterns which reliably generalize to questions whose answers we can’t verify.
Intuitively, ELK aims to directly address the root of the problem: the information asymmetry between a superhuman AI and the humans evaluating its behavior. Our experimental setup aims to capture the scenario where a model is trained on an objective that incentivizes the outputs to not track the truth, despite the model being capable enough to answer truthfully (e.g. a model is mimicking human labeling errors, or has been optimized to make convincing sales).
An ELK method can be useful even when it is no more accurate than other sources of information, as long as its errors are decorrelated with those of existing sources (i.e., model outputs and human annotators).
@@ 2 Related work
Similar to us, Hubinger et al. (2024) finetunes LM’s to exhibit untruthful behavior and MacDiarmid et al. (2024) investigates methods of catching these ”sleeper agents.” However, sleeper agents are models of a more specific type of untruthful behavior called deceptive alignment (Ngo et al., 2024; Hubinger et al., 2019).
@@ 3 Datasets and models
Easy to hard. While the probe or anomaly detector can only be trained reliably on questions that are easy for a human to answer, scalable oversight settings task it with questions hard for a human to answer. However, it is not practical to run experiments in domains where human experts in fact struggle because we would have no source of ground truth against which to measure progress. Bowman et al. (2022) develops the “sandwiching” approach to measuring progress on scalable oversight, in which a scalable oversight method is evaluated against human experts (who are more capable than the model being studied), but is only allowed to make use of non-expert or artificially hindered annotators (who are less capable than the model) for eliciting knowledge from the model. Our settings can be understood in the context of this framework. We discuss our definitions of difficulty in Appendix A.1.
@@
We introduce 12 binary classification datasets, each with an associated example difficulty metric and two label sets: ground truth and systematically untruthful labels. Crucially, the untruthful labels are not generated by merely negating or adding IID noise to ground truth labels, but rather aim to represent somewhat plausible mistakes. For example, in our NLI dataset, untruthful labels correspond to annotation artifacts discovered by Gururangan et al. (2018). Two of our 12 datasets are described below as exemples for the reader, while the remaining 10 are described in Appendix A.
[roundcorner=10pt, backgroundcolor=gray!10] Name: nli
[…]
answer choices: [" Contradict", " Imply"]
Difficulty metric: Pythia eval
Untruthful label: Does the hypothesis contain any negation words (”not”, ”nobody”, ”no”, ”never”, ”nothing”, ”none”)? {mdframed}[roundcorner=10pt, backgroundcolor=gray!10] Name: addition
input: "{op1} + {op2} = {result}. {character}:"
answer choices: [" False", " True"]
Difficulty metric: Magnitude of the smaller operand
We call an example “easy” if it is among the easiest 25% of examples in the dataset, and call it “hard” if it is among the hardest 25%.
@@ 4 Probing methods
Difference-in-means. This method simply sets the weight vector proportional to the difference in mean activations for examples labeled true and false: . The score is the inner product of an activation with the weight vector. Marks and Tegmark (2023) find that it better supports causal interventions on LM activations, and is more robust to distribution shifts.
@@
4.1 Selecting a layer
where is the in-distribution validation AUROC for a probe on layer , is the maximum AUROC over layers, and is the depth of the network. If is empty, we use the middle layer, .
@@ 4.2 Random baseline
4.3 Train and test data
In order to measure generalization of these probing methods, we perform five main transfer experiments, using 4000 train and 1000 test examples, or as many samples as are available for the dataset (see validation and test splits in Fig. 8).
@@ 4.3 Train and test data
4.4 Evaluation: performance gap recovered (PGR)
@@
5 Mechanistic anomaly detection method
As mentioned in the introduction, an alternative approach to ELK is to detect mechanistic anomalies that suggest model output should not be trusted, rather than attempting to elicit reliable output. We fit an anomaly detector without any examples of anomalous behavior, using Alice’s easy examples as a reference distribution. Specifically, we construct a feature vector for each example by concatenating together the probe outputs (log-odds) from all the layers. We then fit a Gaussian to these feature vectors, and use the Mahalanobis distance as the classifier score. For evaluation, we compute the AUROC at distinguishing Bob’s hard examples from Alice’s hard examples.222Following standard practice in the anomaly detection literature, we report AUROC so that we do not have to choose an arbitrary threshold for considering an example to be anomalous. In practice such a threshold would have to be chosen depending on the relative costs of false positives and false negatives. Choosing this threshold may pose challenges in the regime where the anomaly detector is fit on truthful easy examples and we hope for it to classify truthful hard examples as normal too. Notably, both evaluation distributions are different from the reference distribution because in practice one may only have trusted data for easy examples, making it more challenging to identify Bob’s examples. While we focus on the vanilla Mahalanobis distance, Bai et al. (2022) report improved performance using a modified Mahalanobis distance which subtracts the diagonal from the covariance matrix. We experiment with this in Appendix C.
@@ 6 Results
Firstly, the random baseline Fig. 3 demonstrates that reasonably discriminative context-dependent and context-independent representations are somewhat common. Additionally, ground truth knowledge is represented in activations in Bob’s contexts nearly to the same extent as in Alice’s contexts (0.984 PGR) when we allow ourselves to use ground truth supervision to learn it (a logistic regression model trained using Alice’s labels in Bob’s contexts, and evaluated on held out data).
Fig. 4 shows an aggregated summary of our main findings about how probes generalize from truthful to untruthful and from easy to hard examples. The results in Fig. 4(a-b) provide strong evidence for the existence of context-independent representation of knowledge in this setup. However, as seen in Fig. 2, later layers sometimes generalize by mimicking LM output, not context-independently.
@@
Fig. 5 shows a striking property of CCS and CRC: they rarely generalize context-dependently. Their generalization performance is strongly correlated with their in-distribution performance, enabling practitioners to have more confidence that their probe will generalize to untruthful contexts. However, their in-distribution performance is low.
Despite having relatively poor in-distribution AUROC, diff-in-means probes have high AEBH generalization AUROC. Belrose (2023) proves that the difference in means direction has two properties that may help explain this. First, it is guaranteed to have a positive inner product with all discriminative linear probing directions. Second, it is in a certain sense worst-case optimal for additive causal interventions, and prior work has argued that causally explanatory variables are more robust to distribution shifts (Bühlmann, 2018; Schölkopf et al., 2012). In Appendix C.1, we run experiments showing that diff-in-means directions are dramatically more causally implicated in LM output than LogR and LDA.
@@ 8 Conclusions
Our preliminary investigations with this experimental setup demonstrate how it can be of use. First, we find evidence for the existence of context-independent representations of knowledge. Simple probing methods can serve as fairly reliable pointers to truth in contexts where the LM output is untruthful, even on examples harder than any used for training. Our results are informative for scalable supervision of capable systems. For example, if a practitioner finetunes a capable pretrained LM to write complex code, but some fraction of the finetuning data is low-quality and contains security vulnerabilities, then the LM may learn to occasionally write vulnerable code when the prompt resembles low-quality code (Andreas, 2022; McDonell and Reynolds, 2022). Our results suggest the possibility of learning a vulnerability classifier in contexts where we are confident about whether the code has a vulnerability, and that this classifier would generalize to predicting whether the model’s generated code is vulnerability-free in more challenging contexts where the model learned to write insecure code.
@@ Appendix A Datasets
The capitals, hemisphere, and population datasets derive from this Kaggle dataset333https://www.kaggle.com/datasets/viswanathanc/world-cities-datasets?resource=download, which contains information about world cities. Each of these three datasets tasks the LM with verify a basic piece of information about the city.
We also build upon three popular NLP datasets: SciQ [Welbl et al., 2017], amazon polarity [Zhang et al., 2016], and SNLI Bowman et al. [2015]. The untruthful labels for amazon polarity are obtained by checking for the presence of one of 2000 positive sentiment words that can be found here444https://ptrckprry.com/course/ssd/data/positive-words.txt. For these datasets we use the suite of Pythia models to evaluate difficulty. An example’s difficulty is the average cross entropy loss of the Pythia models from scale 160m to 12b, evaluated in a 5-shot setting.
The authors dataset is based on this555Original: https://www.kaggle.com/datasets/jealousleopard/goodreadsbooks
@@ C.1 Causal interventions
Our observations that the difference in means direction is the most causally implicated direction in the LM output corroborate earlier findings by Marks and Tegmark [2023] and Zou et al. [2023], which were later given theoretical grounding in Belrose [2023].
@@ Appendix F Limitations
In practice we likely will not have access to labels about whether an example elicits truthful or untruthful internal mechanisms. One would instead learn supervised probes on arbitrary examples that we can confidently label. Presumably, the LM would also output correct answers on those examples because we can supervise it to do so. Our AEBH experiments therefore aim to capture the scenario where the LM is truthful on examples we can supervise, but not necessarily truthful on examples we can’t supervise. One could also imagine, however, an LM that is always using a mechanism that does not track truth, but that this mechanism only diverges from truth on examples we can’t supervise (e.g. in deceptive alignment; Hubinger et al. [2019], ngo2023alignment). While we do not focus on this, future work could construct datasets and experiments that apply more directly to these scenarios.
@@ Appendix G Future work
We release
666https://github.com/EleutherAI/elk-generalization
Results for this paper can be reproduced from commit 53bad05c4ac52b042f1b0255172f2f425124f670.
Links to the models and data can be found in the README.
our data, models, and code to facilitate reproductions and follow-up work.